OpenStudy (anonymous):

Given f(x)=(3x+5)/(2x-6) find: a) domain, in interval notation b) vertical asymptote(s) if any: c)horizontal asymptote if there is one: thanks your help is very appreciated!

OpenStudy (anonymous):

Do you mean (3x+5)/(2x-6)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a) we know that we can't divide something by 0, hence 2x-6 can't be 0, which means x can't be 3. Therefore, the domain of x is (-infinity, 3) U (3, infinity)

OpenStudy (anonymous):

b) To determine the vertical asymptotes of any given function, we first find at which x value the function does not exist. By looking at the domain we know when x=3, the function does not exist. We now need to test if x=3 is truly an asymptote (which means the limit as x approaches 3 is infinity or -infinity) \[\lim_{x \rightarrow 3+} \frac{ (3x+5) }{ (2x-6) } = \infty\] \[\lim_{x \rightarrow 3-} \frac{ (3x+5) }{ (2x-6) } = -\infty\] x=3 is indeed an asymptote.

OpenStudy (anonymous):

i understand those two....and thank you for taking the time to explain it. i wish i could give you a 100 metals

OpenStudy (anonymous):

are you having trouble explaining it

OpenStudy (anonymous):

c) To see the horizontal asymptote of any given function, we check when the x goes to infinity or -infinity, the y 'stops' at one value. \[\lim_{x \rightarrow -\infty}\frac{ 3x+5 }{ 2x-6 } = \frac{ 3 }{ 2 }\] \[\lim_{x \rightarrow \infty}\frac{ 3x+5 }{ 2x-6 } = \frac{ 3 }{ 2 }\] Hence the horizontal asymptote is y=3/2

OpenStudy (anonymous):

ok that makes sense... thanks a bunch

OpenStudy (anonymous):

np :)